# Bootstrapping Object-level Planning with Large Language Models

### v1

@@ Abstract
We introduce a new method that extracts knowledge from a large language model (LLM) to produce object-level plans, which describe high-level changes to object state, and uses them to bootstrap task and motion planning (TAMP) in a hierarchical manner. Existing works use LLMs to either directly output task plans or to generate goals in representations like PDDL. However, these methods fall short because they either rely on the LLM to do the actual planning or output a hard-to-satisfy goal. Our approach instead extracts knowledge from a LLM in the form of plan schemas as an object-level representation called functional object-oriented networks (FOON), from which we automatically generate PDDL subgoals. Our experiments demonstrate how our method’s performance markedly exceeds alternative planning strategies across several tasks in simulation.
@@ I Introduction
The advent of large language models (LLMs) has led to a plethora of work that exploits their capabilities for a variety of tasks, including planning for robotics [1, 2] and embodied agents [3, 4] using LLMs. Language models encode domain knowledge about the world, which is useful to decision making. These approaches use LLMs as either: 1) a task planner [1, 5, 2], or 2) a task goal generator [6, 7, 8]. As a task planner, a LLM is informed of the task and scene and directly outputs a complete plan, forgoing the use of automated planning [9] with off-the-shelf planners. Plan actions outputted by a LLM are then grounded to action policies or primitives. As a task goal generator, a LLM generates planning definitions in the form of representations like PDDL [10] (short for Planning Domain Definition Language); this type of approach is often situated with task and motion planning (TAMP) [11].
@@
However, existing work in these categories fail to handle complex, goal-oriented tasks in several key aspects. On one hand, positing the LLM as a task planner deprives such methods of guarantees promised by classical planning (viz. optimality and completeness). Recent work has also called to question whether LLMs can effectively plan [12]. On the other hand, using the LLM as a task description generator will fail to generate plan specifications that are guaranteed to work due to the LLM’s lack of embodiment. For instance, it may be difficult for the LLM to generate accurate PDDL definitions simply from a language description of the robot’s environment.
@@
It is natural to explore language models for planning, as they contain domain knowledge useful to planning while often output useful steps. Similarly, LLMs are useful as goal generators because one can still exploit off-the-shelf planners. We propose an approach where the LLM is used to generate partial goal schemas at the object level, which can then be used to generate PDDL subgoals. Such an approach inherits the desirable commonsense planning knowledge of the LLM while still supporting sound and complete task-level planning. The object (as opposed to task) level is the level at which natural language is most appropriate, and at which most knowledge is contained and expressed [13]. While task-level planning focuses on task (action and motion) constraints for task execution, object-level planning focuses on how objects are combined to produce new objects without committing to how these effects will be resolved until runtime.
@@
Figure 1: Our approach prompts a LLM for object-level information, from which we construct an object-level plan (as FOON). This plan schema bootstraps task- and motion-level planning (TAMP) with PDDL subgoals.
@@
We propose a modularized approach that distills domain knowledge from a LLM to generate object-level plans [14], which can then be used to bootstrap hierarchical planning. We situate object-level planning as an interface between human language and TAMP and exploit an object-level representation (OLR) called the functional object-oriented network (FOON) [15]. Recent work has shown how object-level knowledge in FOON can automatically generate PDDL subgoals [16]; however, this assumes that partial plan specifications already exist as a FOON. In this work, we exploit the capabilities of LLMs for generating FOONs which, when coupled with LLMs, overcomes the inability of LLMs to directly output feasible task plans while exploiting the higher, object-level nature of LLM output and language on a whole.
@@
Figure 2: Our approach interfaces with a language model to generate object-level plans (as FOON graphs) for bootstrapping task and motion planning. We generate task-level subgoals as PDDL subgoal definitions by grounding object-level subgoals to a robot’s environment; with these task-level definitions, we perform task planning to obtain task plan segments per object-level action, which we execute using motion-level planning, improving upon prior work [16].
@@
The contributions of our work are as follows: 1) we introduce a modularized planning approach (Figure 1) that interfaces with a LLM to generate natural language instructions, from which we transform into an OLR (e.g., FOON) for hierarchical planning; 2) we show how object-level information can be distilled directly from an LLM and then used to generate planning definitions as PDDL, improving the feasibility of generated plans; and 3) we showcase markedly better performance than alternative LLM-based methods.
@@ II Background
Large Language Models: A large language model (LLM) is a complex neural network model trained via self-supervised learning and self-attention (the latter revolutionized by the transformer architecture [17]). LLMs have shown remarkable performance in natural language processing (NLP) and generation tasks. LLM variants such as GPT [18] and LLaMA [19] are trained on large corpora of text gathered from the internet and fine-tuned using RLHF (reinforcement learning from human feedback). For this reason, a LLM can be thought of as a “compressed” representation of domain knowledge from the web [20], which is why we aim to exploit these models to inform planning. This work uses OpenAI’s Chat-GPT [21].
Task and Motion Planning: The aim of task and motion planning (TAMP) is to integrate higher-level symbolic task planning with lower-level motion planning for enabling robots to solve complex long-horizon tasks [11]. At the lower level, motion planning finds collision-free robot motions or trajectories that will resolve a task. However, typical robot tasks are too complex to be solved via motion-level planning alone due to the high-dimensional nature of the robot’s task and configuration space. For this reason, task planning is necessary as an added layer to reason over an abstraction of a robot’s actions and its environment for simplifying motion planning. Task planning assumes a state description $\mathcal{S}$ using logical predicates, which are true or false depending on whether or not the robot observes them. Starting from an initial state $s_{0}\in\mathcal{S}$, task planning finds an action sequence $a\in\mathcal{A}$ that achieves a goal $g$ as a task plan $\mathcal{P}=\{a_{1},...,a_{n}\}$. An action $a$ refers to a robot-executable skill or policy; our work assumes access to a repertoire of skills, which we denote by $\mathcal{A}$, that are defined as planning operators in PDDL [10].
@@ III Object-level Planning with Language Models
There exists a disconnect between language and TAMP, which makes it unsuitable for generalization across tasks and settings. Yet, existing works use LLMs either as planners or task description generators for task execution; these approaches fall short due to their overestimation of a LLM’s ability to reason about task- and motion-level constraints. It is impractical to provide the entire context of a task setting to a LLM and expect it to handle all the reasoning about a robot’s embodiment (e.g., where objects are located, in what poses they are, what type of gripper the robot has, etc.) in order to generate adequate planning definitions or feasible task plans.
Instead, the strong suit of language models lies in its ability to provide plan sketches or subgoals that are useful to decision making at both task and motion levels. This is because language models can express task-relevant knowledge in a generic yet informative way using natural language. Imagine your typical cooking recipe, for instance: a recipe provides a sketch of object interactions agnostic to the state of your kitchen or the recipe writer’s kitchen. It also does not provide details on how actions should be executed (e.g., which hand should you use, how should you grasp an object, etc.). What a recipe expressed in natural language may provide however is an idea of the types of actions and object-object interactions are required to complete a task. Rather, the exact details of task- and motion-level execution are resolved at run time.
It is for this reason why we adopt an object-level planning approach to bootstrap task and motion planning. We generate object-level plan sketches, which provide task-level subgoals that naturally interface language and decision making, using a LLM. Briefly, given a prompt for a task to be executed by a robot, our approach (Figure 2) uses a LLM to generate a sequence of natural language instructions, which is then transformed into an object-level plan (OLP) represented as a FOON. It is then through task planning where properties relevant to the robot, e.g., robot’s end-effector and object poses) are used to find a task-aware plan which we then execute via TAMP. We perform task-level planning by transforming each OLP action into PDDL definitions to find task plan segments.
@@ III-A Object-level Planning
In this work, we add another level of planning situated above TAMP called object-level planning, which reasons solely over object interactions [14]. We situate object-level planning with the functional object-oriented network (FOON): a knowledge graph representation describing object-action relations [15]. Formally, a FOON $\mathcal{G}=\{\mathcal{O},\mathcal{M},\mathcal{E}\}$ is a bipartite graph with object nodes ($o\in\mathcal{O}$) and motion nodes ($m\in\mathcal{M}$) connected via directed edges ($e\in\mathcal{E}$), which reflect the change of an object’s state as it is manipulated via a corresponding action. An object $o=(o_{t},o_{s},o_{\mathcal{I}})$ is defined as a tuple with the following attributes: its object type or name ($o_{t}$), its state ($o_{s}$), and, if applicable, its object composition ($o_{\mathcal{I}}=\{o_{t_{1}},o_{t_{2}},...,o_{t_{n}}\}$, where $n=|\mathcal{I}|$). A motion node $m=(m_{t})$ is defined by an action verb or type ($m_{t}$). A FOON describes object-state transitions via functional units ($\mathcal{FU}=\{\mathcal{O}_{in},\mathcal{O}_{out},\tilde{m}\}$) at a level closer to human language. A functional unit defines preconditions and effects of executing an action ($\tilde{m}$), where a set of input nodes ($\mathcal{O}_{in}$) are required to produce a new set of output object nodes ($\mathcal{O}_{out}$). We illustrate an example of a functional unit as Figure 3, which describes an action for a block stacking task. Foundation models naturally interface with object-level representations like FOON due to its similarity to human language, which in turn allows it to interface with tools that combine vision and language [18, 22].
@@
III-B LLM Prompting to Object-level Plan
Our goal is to extract attributes for object-state transitions before and after each action is performed (i.e., preconditions and effects) to construct an OLP describing task subgoals. We instantiate object-level planning with FOONs. Our approach constructs a FOON $\mathcal{G}_{\mathcal{T}}$, where $\mathcal{T}$ is a task given in natural language (such as “Make a tower of two red blocks”—see Figure 3) via a two-stage process. In addition to the task instruction $\mathcal{T}$, we include a language description of objects in the scene (from which the LLM will determine which ones are relevant to the task) as well as a set of example object-level plans (as FOONs) $\mathcal{X}_{\mathcal{G}}=\{\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{n}\}$ for reference.
The first stage involves prompting a LLM for a plan sketch of natural language instructions denoted by $\mathcal{P}_{L}=\{\xi_{1},\xi_{2},...,\xi_{n}\}$, where $\xi_{i}$ refers to an instruction as text. As an example in Figure 3, given a task and available objects (without any context about their present configuration), we expect text instructions $\mathcal{P}_{L}$ that solely mention red blocks for the task “Make a tower of 2 red blocks.” During this step, we transform the top-$k$ most similar FOON in $\mathcal{X}_{\mathcal{G}}$ into example plan sketches, from which the LLM selects the one closest to the new task to use as reference (denoted by $\hat{\mathcal{G}}$). We identify the top $k$ candidates using cosine similarity between text embeddings of the task prompt $\mathcal{T}$ and the set of instructions for a given reference $\hat{\mathcal{G}}\in\mathcal{X}_{\mathcal{G}}$. An example sketch may describe how three generic blocks (regardless of type) can be stacked into a tower. In the second stage, the LLM must reason about each instruction $\xi_{i}\in\mathcal{P}_{L}$ to generate an OLP for the novel task. We prompt the LLM to reason about state changes of task-relevant objects, specifically geometric relations for task-level planning (Section IV-A). In the previous example, we expect output with state descriptions such as “first red block on second red block,” “first red block under nothing,” “second red block under first red block,” and “second red block on table” (Figure 3). We assist the LLM by providing $\hat{\mathcal{G}}$ in the prompt, with which it must generate a new FOON $\mathcal{G}_{\mathcal{T}}$ for the novel task. Inspired by prior work on code writing for robots [23], we codify $\hat{\mathcal{G}}$ as a JSON. The LLM then outputs a codified OLP $\tilde{\mathcal{G}}_{\mathcal{T}}$, capturing each instruction $\xi_{i}\in\mathcal{P}_{L}$. Finally, each action in $\tilde{\mathcal{G}}_{\mathcal{T}}$ forms a functional unit $\mathcal{FU}_{i}\in\mathcal{G}_{\mathcal{T}}$.
IV Bridging to Task and Motion Planning
We generate a plan schema $\mathcal{G}_{\mathcal{T}}$, with which we can solve a task $\mathcal{T}$ given in natural language. However, this schema is too abstract to be executed in its present form, and it must be grounded to the robot’s embodiment and environment [16]. Therefore, we use $\mathcal{G}_{\mathcal{T}}$ to bootstrap TAMP by providing PDDL subgoals. This is done through a hierarchical approach that automatically transforms each object-level action in $\mathcal{G}_{\mathcal{T}}$ into PDDL problem definitions and searches for a robot-executable plan given a predefined set of robot skills or operators.

### v3

@@ Abstract
We introduce a new method that extracts knowledge from a large language model (LLM) to produce object-level plans, which describe high-level changes to object state, and uses them to bootstrap task and motion planning (TAMP). Existing work uses LLMs to directly output task plans or generate goals in representations like PDDL. However, these methods fall short because they rely on the LLM to do the actual planning or output a hard-to-satisfy goal. Our approach instead extracts knowledge from an LLM in the form of plan schemas as an object-level representation called functional object-oriented networks (FOON), from which we automatically generate PDDL subgoals. Our method markedly outperforms alternative planning strategies in completing several pick-and-place tasks in simulation. [^2] Project Website: https://davidpaulius.github.io/olp_llm/
@@ I Introduction
The advent of large language models (LLMs) has led to a plethora of work that exploits their capabilities for a variety of tasks, including planning for robotics [1, 2] and embodied agents [3, 4]. These approaches use LLMs as either a planner [1, 5, 2], or a goal generator [6, 7, 8, 9]. As a task planner, an LLM is informed of the task and scene and directly outputs a complete plan, thus forgoing automated planning with off-the-shelf planners [10]. Plan actions generated by an LLM are then grounded to action policies or primitives. As a task goal generator, an LLM generates planning definitions in the form of representations like PDDL [11] (short for Planning Domain Definition Language); this type of approach is often associated with task and motion planning (TAMP) [12].
@@
However, existing work in these categories fails to handle complex, goal-oriented tasks in several key aspects. On the one hand, positing the LLM as a task planner deprives such methods of guarantees promised by classical planning (viz. optimality and completeness). Recent work has also called to question whether LLMs can effectively plan [13]. On the other hand, using the LLM as a task description generator will fail to generate plan specifications that are guaranteed to work due to the LLM’s lack of embodiment. For instance, it may be difficult for the LLM to generate accurate PDDL definitions simply from a language description of the robot’s environment.
@@
It is natural to exploit language models for planning as they contain useful domain knowledge and often output useful steps. Similarly, they are useful as goal generators because one can still rely on off-the-shelf planners. This work uses an LLM to generate partial goal schemas at the object level, which can then form PDDL subgoals. Such an approach inherits the desirable commonsense planning knowledge of the LLM while supporting sound and complete task-level planning. The object (as opposed to task) level is the level at which natural language is most appropriate and at which most knowledge is captured and expressed [14, 15]. While task-level planning focuses on action or motion constraints for execution, object-level planning focuses on object interactions without committing to how these effects will be resolved until runtime.
@@
Figure 1: Our approach prompts an LLM for object-level information with which we construct an object-level plan (as FOON). This plan schema bootstraps task- and motion-level planning (TAMP) via PDDL subgoals.
@@
We propose a modular approach that distills domain knowledge from an LLM to generate object-level plans [15], which then bootstrap hierarchical planning. We situate object-level planning as an interface between human language and TAMP and exploit an object-level representation (OLR) called the functional object-oriented network (FOON) [16]. Recent work has shown how object-level knowledge in FOON can automatically generate PDDL subgoals [17]; however, this assumes that partial plan specifications already exist as a FOON. We exploit the capabilities of LLMs for object-level planning, overcoming the inability of LLMs to directly output feasible task plans while exploiting the higher, object-level nature of LLM output and language as a whole.
@@
Figure 2: Our approach interfaces with a language model to generate object-level plans (as FOON graphs) for bootstrapping task and motion planning. We generate task-level subgoals as PDDL subgoals by grounding object-level subgoals to the robot’s environment; with these task-level definitions, task planning to obtains task plan segments per object-level action, which are executed using motion-level planning, improving prior work [17].
@@
The contributions of our work are as follows: first, we introduce a modular planning approach (Figure 1) that interfaces with an LLM to generate natural language instructions, from which we transform into an OLR (e.g., FOON) for hierarchical planning. Second, we show how object-level information can be distilled directly from an LLM and then used to generate planning definitions as PDDL, improving the feasibility of generated plans. Finally, we showcase markedly better performance than alternative LLM-based methods.
@@ II Background
Large Language Models: A large language model (LLM) is a complex neural network model trained via self-supervised learning and self-attention [18]. LLMs have shown remarkable performance in natural language processing (NLP) and text generation tasks. Variants such as GPT [19] and LLaMA [20] are trained on large corpora of text collected from the Internet and fine-tuned using RLHF (reinforcement learning from human feedback). For this reason, an LLM can be thought of as a “compressed” representation of domain knowledge from the web [21], which is why we aim to exploit these models to inform planning. This work uses OpenAI’s Chat-GPT [22].
Task and Motion Planning: The aim of task and motion planning (TAMP) is to integrate higher-level symbolic task planning with lower-level motion planning to enable robots to solve complex long-horizon tasks [12]. At the lower level, motion planning finds collision-free robot motion or trajectories that are typically used to achieve a task. However, typical robot tasks are too complex for motion-level planning alone. For this reason, task planning is necessary as an added layer to reason over an abstraction of the robot’s actions and environment. Task planning assumes a state description $\mathcal{S}$ using logical predicates, which are true or false depending on whether or not the robot observes them. Starting from an initial state $s_{0}\in\mathcal{S}$, task planning finds an action sequence $a\in\mathcal{A}$ that achieves a goal $g$ as a task plan $\mathcal{P}=\{a_{1},...,a_{n}\}$ [10]. An action $a$ refers to a robot-executable skill or policy; our work assumes access to a repertoire of skills, which we denote by $\mathcal{A}$, which are defined as planning operators in PDDL [11]. Finally, given the task plan $\mathcal{P}$, motion planning finds collision-free trajectories that reproduce the intended effects of each action; our work uses OMPL [23] for motion planning.
II-A Related Work
Language Models for Planning: Many researchers have explored the use of language models for robotics applications, having been inspired by their remarkable performance in language-related tasks. Prior works have investigated the planning capabilities of LLMs [24, 13]. Other works supplement task planning with language models [6, 25, 5, 26, 8, 27]. LLM+P [6] generates PDDL problem file via LLM prompting. Much like our work, existing works use LLMs as an informer of subgoals for classical planning [26, 8, 27, 9]. In particular, DELTA [8] resembles our method in that it decomposes a task into a series of PDDL subgoal definitions directly output by an LLM. Our approach uses an LLM at the object level and not task level (i.e., PDDL). Recent work also iteratively prompts an LLM for FOON generation [28]. Similar to DELTA, they do not focus on generating nor executing physically valid plans.
Language Models as Planners: Several works treat language models as robotic task planners. SayCan [1] combines a language model and affordance detectors for driving robotic execution given a task prompt. PaLM-E [2] is an embodied language model that directly incorporates continuous observations (like images, state estimates, or other sensor modalities) into the language embedding space. These works have shown that language models are capable of performing some degree of embodied reasoning. However, one major drawback of these works is that they require a large amount of engineering effort, particularly to enable them to operate in novel environments and solve long-horizon tasks. Previous works also exploited the reasoning capabilities of an LLM to solve a wide range of tasks both in simulation [3, 29, 4] and with a real robot [4].
@@ III Object-level Planning with Language Models
There exists a disconnect between language and the task level, which makes TAMP unsuitable for generalization across tasks and settings. Yet, existing works use LLMs either as planners or task description generators for task execution; these approaches fall short because of the inability of LLMs to correctly reason about task- and motion-level constraints. It is impractical to provide the entire context of a task setting to an LLM and expect it to handle all the reasoning about a robot’s embodiment (e.g., where objects are located, in what poses they are, what type of gripper the robot has, etc.) in order to generate adequate planning definitions or feasible task plans.
Instead, the strength of language models lies in their ability to provide approximate subgoals that are useful to decision making at both task and motion levels. This is because language models can express task-relevant knowledge in a generic yet informative way. Imagine your typical cooking recipe, for instance: a recipe provides a sketch of object interactions agnostic to the state of the reader’s kitchen or the recipe writer’s kitchen. It also does not provide details on how actions should be executed (e.g., which hand should be used, how should an object be grasped, etc.). What a recipe expressed in natural language may provide, however, is an idea of the types of actions and inter-object interactions necessary to complete a task. Rather, the exact details of task- and motion-level execution are resolved at run time.
For these reasons, we adopt an object-level planning approach to bootstrap task and motion planning [17]. We generate object-level plan sketches, which provide task-level subgoals that naturally interface language and decision making, using an LLM. Briefly, given a language command to a robot, our approach (Figure 2) uses an LLM to generate a sequence of natural language instructions, which is then transformed into an object-level plan (OLP) represented as a FOON. It is then through task planning where properties relevant to the robot (e.g., robot’s end-effector and object poses) are used to find a task-aware plan which is then executed via TAMP. Task-level planning is achieved by transforming each OLP action into PDDL definitions to find task plan segments.
@@ III-A Object-level Planning
We adopt another layer of planning above TAMP called object-level planning, which considers changes to object state [15]. We use an object-level representation in the form of a knowledge graph called the functional object-oriented network (FOON) [16, 30]. Formally, a FOON $\mathcal{G}=\{\mathcal{O},\mathcal{M},\mathcal{E}\}$ is a bipartite graph with object nodes ($o\in\mathcal{O}$) and motion nodes ($m\in\mathcal{M}$) connected via directed edges ($e\in\mathcal{E}$), which reflect the change of an object’s state as it is manipulated via a corresponding action. An object $o=(o_{t},o_{s},o_{\mathcal{I}})$ is defined as a tuple with the following attributes: its object type or name ($o_{t}$), its state ($o_{s}$), and, if applicable, its object composition ($o_{\mathcal{I}}=\{o_{t_{1}},o_{t_{2}},...,o_{t_{n}}\}$, where $n=|\mathcal{I}|$). A motion node $m=(m_{t})$ is defined by an action verb or type ($m_{t}$). A FOON describes object-state transitions via functional units ($\mathcal{FU}=\{\mathcal{O}_{in},\mathcal{O}_{out},\tilde{m}\}$) at a level close to human language. A functional unit defines preconditions and effects of executing an action ($\tilde{m}$), where a set of input nodes ($\mathcal{O}_{in}$) are required to produce a new set of output object nodes ($\mathcal{O}_{out}$). We illustrate an example of a functional unit in Figure 3, which describes an action for a block stacking task. We argue that foundation models naturally interface with object-level representations due to their similarity to human language, which in turn allows them to interface with tools that combine vision and language [19, 31].
III-B LLM Prompting to Object-level Plan
Our goal is to extract attributes for object-state transitions before and after each action is performed (i.e., preconditions and effects) to construct an OLP describing task subgoals. We instantiate object-level planning with FOONs. Our approach constructs a FOON $\mathcal{G}_{\mathcal{T}}$, where $\mathcal{T}$ is a task given in natural language (such as “Make a tower of two red blocks”—see Figure 3) via a two-stage process. In addition to the task instruction $\mathcal{T}$, we include a language description of objects in the scene (from which the LLM will determine those that are relevant to the task) as well as a set of example object-level plans (as FOONs) $\mathcal{X}_{\mathcal{G}}=\{\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{n}\}$ for reference.
The first stage prompts an LLM for a plan sketch comprised of natural language instructions denoted by $\mathcal{P}_{L}=\{\xi_{1},\xi_{2},...,\xi_{n}\}$, where $\xi_{i}$ refers to an instruction as text. As an example in Figure 3, given a task and available objects (without any context about their present configuration), we expect text instructions $\mathcal{P}_{L}$ that solely mention red blocks for the task “Make a tower of 2 red blocks.” During this step, we transform the top-$k$ most similar FOONs in $\mathcal{X}_{\mathcal{G}}$ into example plan sketches, from which the LLM selects the one closest to the new task to use as reference (denoted by $\hat{\mathcal{G}}$). We identify the top $k$ candidates using cosine similarity between text embeddings of the task prompt $\mathcal{T}$ and the set of instructions for a given reference $\hat{\mathcal{G}}\in\mathcal{X}_{\mathcal{G}}$. An example sketch may describe how three generic blocks (regardless of type) can be stacked into a tower. In the second stage, the LLM must reason about each instruction $\xi_{i}\in\mathcal{P}_{L}$ to generate an OLP for the novel task. We prompt the LLM to reason about state changes of task-relevant objects, specifically geometric relations for task-level planning (Section IV-A). In the previous example, we expect output with state descriptions such as “first red block on second red block,” “first red block under nothing,” “second red block under first red block,” and “second red block on table” (Figure 3). We assist the LLM by providing $\hat{\mathcal{G}}$ in the prompt, with which it must generate a new FOON $\mathcal{G}_{\mathcal{T}}$ for the novel task. Inspired by previous work on code writing for robots [32], we codify $\hat{\mathcal{G}}$ as a JSON. The LLM then outputs a codified OLP $\tilde{\mathcal{G}}_{\mathcal{T}}$, which captures each instruction $\xi_{i}\in\mathcal{P}_{L}$. Finally, each action in $\tilde{\mathcal{G}}_{\mathcal{T}}$ forms a functional unit $\mathcal{FU}_{i}\in\mathcal{G}_{\mathcal{T}}$.
IV Bridging to Task and Motion Planning
We generate a plan schema $\mathcal{G}_{\mathcal{T}}$, with which we can solve a task $\mathcal{T}$ given in natural language. However, this schema is too abstract to be executed in its present form, and it must be grounded to the robot’s embodiment and environment [17]. Therefore, we use $\mathcal{G}_{\mathcal{T}}$ to bootstrap TAMP via PDDL subgoals. This is done through a hierarchical approach that automatically transforms each object-level action in $\mathcal{G}_{\mathcal{T}}$ into PDDL problem definitions and searches for a robot-executable plan given a predefined set of robot skills or operators [17].